A coin is tossed until a head appears, or the tail appears 4 times in succession, then the probability distribution of the number of tosses is
Answer & explanation
Correct answer: option 1
The correct answer is Option (1) → (table in Option 1)
Let X be the number of tosses. Tossing stops at the first head, or after 4 tosses in any case.
X = 1: H, so P = 1/2
X = 2: TH, so P = 1/4
X = 3: TTH, so P = 1/8
X = 4: TTTH or TTTT, so P = 1/16 + 1/16 = 1/8
Check: 1/2 + 1/4 + 1/8 + 1/8 = 1.